For a certain time period the ratio of the dollar value of exports to the dollar value of imports for the United States was 2 to 3. If the value of exports during that time period was 48 billion dollars, then what was the value of imports?
Answer by edjones(8007)

x/48 = 3/2
x=48*3/2
=72 billion dollars.
.
Ed